OpenStudy (anonymous):

Find two positive numbers such that the sum of the first and twice the second is 120 and whose product is a maximum.

OpenStudy (anonymous):

x + 2y = 120 2xy is a maximum combinig P = 2xy = x(120 - x) differentiate dP/dx = 120 - 2x = 0 for maxm/minm so x = 60 second derivative = -2 so this gives maximum 60 + 2y = 120 so y = 30 your 2 positive numbers are 60 and 30

OpenStudy (anonymous):

Thanks! Can you help me with this one too?... Find the length and width of a rectangle that has an area of 968 square feet and whose perimeter is a minimum.

OpenStudy (anonymous):

sorry guy - its very late here in UK and i gotta get up early

OpenStudy (anonymous):

It's okayy